# Dynamic and Adaptive Feature Generation with LLM

### v1

@@ 3.3 Feedback and Optimization
We now have a new feature set composed of all $l$ subsets, denoted as $\mathcal{F}_{l}^{{}^{\prime}}={\{\mathcal{F}_{l},g_{1},\ldots,g_{k}\}}$. After the agents generate all $l$ subsets, we move forward to evaluate the performance of these subsets in downstream tasks and the effectiveness of the generated nodes at this layer. We then gather the feedback of downstream tasks by the performance metrics $\mathbf{\theta}=\{\theta_{t}\}$, for each iteration $t$ corresponding to each generation layer. This feedback reflects the effectiveness of the newly created feature subset’s effectiveness and illustrates the agents’ contribution to the task. The agents evaluate their performance by comparing the metric $\theta_{t}$ with $\theta_{(t-1)}$. If $\theta_{t}$ significantly improves over $\theta_{(t-1)}$, the feature generation strategy is considered “effective”. The agents’ self-evaluation positively correlates with the increase in these metrics; a larger improvement signifies a more successful feature generation strategy.
Furthermore, each agent shares their thought process and inferences when generating features, including the logic behind selecting specific operations. This transparency helps agents understand and adopt diverse strategies to bring collaborative improvement. Through self-evaluation and peer interactions, agents identify which strategies successfully enhanced performance and which may require adjustments or replacements. In the next round, agents generate increasingly effective feature subsets adopting $\mathcal{F}_{l}^{{}^{\prime}}$ and $\mathcal{O}$ to form the next layer of generated nodes. The iterative process continues until the optimal feature set $\hat{\mathcal{F}}$ is found or the maximum number of iterations $T$ is reached:
@@
Monte Carlo Tree Search. In order to further optimize the feature space and balance exploration and exploitation, we employ an improved Monte Carlo Tree Search (MCTS) to explore new or underutilized feature space structures after $T$ rounds of iterations. We adopt this MCTS framework because of its random sampling-based search strategy, which enhances the probability of finding superior feature combinations. In our case, this flexible searching method builds the tree on all agents’ generation nodes and adjusts the search direction based on the performance feedback. Here, we define the root node as the original feature set given to the LLM, with each generation node representing a tree node. The search covers all generation nodes in the last iteration layer $Layer_{T}$ as leaf nodes.
@@
We start with node evaluation by calculating the improvement in the downstream task’s performance compared to the parent node. The evaluation reflects the effectiveness of newly generated features. Then, the MCTS iteratively builds a search tree by cycling through four phases after evaluation: selection, expansion, evaluation, and searching. In the selection phase, children nodes are recursively selected from the root utilizing the Upper Confidence Bound (UCB) (Auer et al., 2002):
@@
where $s_{i}$ and $s^{{}^{\prime}}_{i}$ are the visit counts for node $i$ and its parent respectively, $C$ is a hyperparameter balancing exploration and exploitation. The value $w_{i}$ represents the average performance improvement of all descendant nodes of node $i$, computed based on changes in the downstream task performance metric $\mathbf{\theta}$. Specifically, for each node $i$, the value of $w_{i}$ is calculated as the average effect of performance improvements across all its children in the downstream tasks, given by:
$w_{i}=\frac{1}{n}\sum_{j=1}^{n}(\theta_{t}^{(j)}-\theta_{(t-1)}^{(j)}).$ (6)
Here, $\theta_{t}^{(j)}$ and $\theta_{(t-1)}^{(j)}$ represent the performance metrics at consecutive time points $t$ and $t-1$ for node $j$, indicating the performance of node $j$, and $n$ is the total number of nodes considered.
In the tree, we select the most potential node $s_{l}$. We define the selection criterion for each node as:
@@
$\mathcal{F}_{l}^{\prime}=\mathcal{O}(\mathcal{F}_{l}^{\text{select}}),$ (8)
where $\mathcal{O}$ represents the set of operations applied to transform $\mathcal{F}_{l}^{\text{select}}$ into $\mathcal{F}_{l}^{{}^{\prime}}$. At the end of the MCTS process, the optimal feature subset is determined by choosing the node with the highest value of $w_{i}$ from among the leaf nodes of the tree, representing the feature set that maximized performance improvement during the simulations.
$\mathcal{F}^{*}=\max_{\mathcal{F}_{l}\in\text{Leaf nodes}}w_{l}.$ (9)
This $\mathcal{F}^{*}$, represents the optimal feature subset identified by the MCTS, which should then be validated further in real-world tasks to confirm its efficacy.
@@ 4 Experiments
In this section, we present three experiments to demonstrate the effectiveness of the LFG. First, we compare the LFG against several baseline methods on multiple downstream classification tasks. Second, we perform a robustness check on the performance improvement of LFG. Finally, we further study and analyze the iterative performance improvements in the experimental results and discuss the reason.
@@ 4.1 Experimental Setup
Datasets. We evaluate the LFG method on four real-world datasets from UCI, including Ionosphere (Ino Sigillito et al. (1989), Amazon Commerce Reviews (Ama) Liu (2011), and Abalone (Aba)  Nash et al. (1995), as well as Diabetes Health Indicators Dataset (Dia) Teboul (2022) from Kaggle. The detailed information is shown in Table 1. For each dataset, we randomly selected $55\%$ of data for training.
Table 1: Datasets Description.
@@
Metrics. We evaluate the model performance by the following metrics: Overall Accuracy (Acc) measures the proportion of true results (both true positives and true negatives) in the total dataset. Precision (Prec) reflects the ratio of true positive predictions to all positive predictions for each class. Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class. F-Measure (F1) is the harmonic mean of precision and recall, providing a single score that balances both metrics.
Downstream Tasks. We apply the LFG model across a range of classification tasks, including Random Forests (RF), Decision Tree (DT), K-Nearest Neighbor (KNN) and Multilayer Perceptrons (MLP). We compare the performance outcomes in these tasks both with and without our method.
Baseline Models. We compare the LFG method with raw data (Raw), the Least Absolute Shrinkage and Selection Operator (Lasso), and Feature Engineering for Predictive Modeling using Reinforcement Learning Khurana et al. (2018) (RL). Here, we set the same operation set for both LFG and RL method to consist of square root, square, cosine, sine, tangent, exp, cube, log, reciprocal, sigmoid, plus, subtract, multiply, and divide. For LLM, we perform all the experiments on OpenAI API, GPT-3.5 Turbo model through the OpenAI platform  [^1] https://openai.com/index/openai-api/.
@@ 4.2 Experimental Results
Overall Performance. Table 2 shows the overall results of LFG and the baseline models’ performance. From the table we can see that the LFG method consistently surpasses baseline methods across a variety of metrics and datasets. Specifically, we show the results of LFG in 3 iterations of generation, denoted as LFG-3, to compare with LFG in $T\leq 10$ (LFG). On dataset Ion the LFG achieves the best accuracy of $95.6\%$, which is $6.4\%$ increase over raw data on task RF, and $4.5\%$ over RL method. On dataset Ama, the LFG-3 improves the accuracy of raw data of $59.6\%$ to $61.5\%$ while LFG performs a increase of $4.3\%$ on RF. In precision, the LFG also shows superior performance. For example, on dataset Ion the LFG reaches a precision of $90.6\%$, surpassing the highest precision of $86.7\%$ among all baselines on KNN. This result indicates that the LFG is effective in reducing misclassifications. Regarding the metric recall, the LFG increases the Ion’s recall by $9.8\%$ comparing with vanilla data and by $4.3\%$ comparing with best of baselines on DT task, which demonstrates that LFG brings higher improvement on identifying positive samples. For F1, LFG on Ama improves the data performance from $63.8\%$ to $67.6\%$ in three iterations, and the F1 score increases up tp $68.0\%$. As F1 Score is the harmonic mean of precision and recall, this improvement shows the LFG’s capability to balance between reducing misclassifications and minimizing missed classifications.
@@
Robustness Check. We validate the experiments through five-fold cross-validation. The LFG methods demonstrate robust performance across various datasets and achieve notable metric improvements, including $2.72\%$ in accuracy and approximately $5\%$ in precision, recall and F1 scores with LFG-3, and further $4.81\%$ in accuracy and approximately $7\%$ in other metrics with LFG. The results illustrate the adaptability and effectiveness of the LFG. The extended iterations continue to refine and optimize the feature space, which is reflected in the continuous improvements in precision and recall. The consistent enhancement across all metrics demonstrates the LFG’s reliability and stability comprehensively.
Iterative Performance Improvements. We move forward comparing the performance improvements of LFG-3 and LFG. On one hand, the incremental improvement from LFG-3 to LFG illustrates the continuous enhancement in all metrics, indicating that during additional iterations, the model refines and adjusts its generation strategy to explore better solutions for the feature space. On the other hand, the results of LFG-3 with only three iterations already shows preservative enhancement in all datasets. This shows that the model effectively and rapidly finds the zones of optimal feature space within limited steps of searching. Showing in Figure 5, we can see that the average number of features in a feature subset increases continuously in the first $5$ rounds of generation. The increase in the number of features implies that the the feature generation algorithm is actively finding new, potentially feature combinations in the feature space and is capable of expanding the feature space effectively.
@@ 5 Conclusion
In this paper we present a novel feature generation method LFG utilizing LLM to enhance the interpretability, adaptability, and strategic flexibility of automated feature engineering. Our target in designing the model is to address the challenges in traditional feature generation including lack of explainability, limited applicability, and rigid strategy formulation with expert-level LLM agents. These problems can significantly limit the broader deployment of feature engineering on machine learning tasks in diverse scenarios. Thus we present LFG with a dynamic and adaptive automated feature generation approach to enhance the interpretability and extends the utility across various data types and tasks. Moreover, we present extensive experimental results to demonstrate that our approach outperforms various existing methods. For future work, we plan to extend this automated feature engineering paradigm across various feature engineering techniques and machine learning tasks, and to conduct a thorough investigation into the transformations that our approach applies to the feature space to study deeper insights into its underlying mechanisms and impacts.
@@ 6 Limitations and Ethics Statements
While our LFG shows significant advancements and wide adaptability, there are several limitations that require further exploration, including high computational demands and limited scalability with very large or complex datasets. Second, the effectiveness of the generated features heavily relies on the quality of input data, which can affect the model’s performance in scenarios with poorly curated or noisy datasets. Finally, there may be inherent limitations of adopting LFG in tasks with unique requirements.
@@
The LFG framework can enhance the effectiveness and explainability of the feature generation process with LLM. However, as the framework uses pre-trained GPT-3.5 Turbo as the generation model, it may inherit the ethical concerns associated with GPT-3.5 Turbo, such as responding to harmful queries or exhibiting biased behaviors.
Appendix A Appendix
A.1 Definitions and Symbols
Table 3: Mathematical Symbols in Feature Set and Operation Set.
Symbol Explanation
$\mathcal{F}$ Feature set, represented as =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$, where $f_{i}$ is the $i$-th feature.
$\mathcal{D}$ Dataset, represented as a series of samples and their corresponding labels, $\mathcal{D}=\langle\mathcal{F},\mathcal{y}\rangle$.
$\mathcal{y}$ Target label set, corresponding to the samples in the dataset $\mathcal{D}$.
$\mathcal{O}$ Operation set, containing all available mathematical operations, $\mathcal{O}=\mathcal{O}_{\text{unary}}\cup\mathcal{O}_{\text{binary}}$.
$\mathcal{O}_{\text{unary}}$ Unary operation set, e.g., $\mathcal{O}_{\text{unary}}=\{f_{1}^{2},e^{f_{1}},\log(f_{1}),\ldots\}$.
$\mathcal{O}_{\text{binary}}$ Binary operation set, e.g., $\mathcal{O}_{\text{binary}}=\{f_{1}+f_{2},f_{1}\times f_{2},f_{1}/f_{2},\ldots\}$.
$\mathcal{R}$ Downstream machine learning task (e.g., classification, regression, ranking, detection).
$\phi$ Performance indicator of the downstream task $\mathcal{R}$.
$\hat{\mathcal{F}}$ Updated feature set, derived by applying operations from $\mathcal{O}$ to the original feature set $\mathcal{F}$.
A.2 Definitions of Feature Operations
[…]
    $\text{square}(f_{i})=f_{i}^{2}$ (10)
[…]
    $\text{exp}(f_{i})=e^{f_{i}}$ (11)
[…]
    $\text{log}(f_{i})=\log(f_{i})$ (12)
[…]
    $\text{sin}(f_{i})=\sin(f_{i})$ (13)
[…]
    $\text{cos}(f_{i})=\cos(f_{i})$ (14)
[…]
    $\text{plus}(f_{i},f_{j})=f_{i}+f_{j}$ (15)
[…]
    $\text{multiply}(f_{i},f_{j})=f_{i}\times f_{j}$ (16)
[…]
    divide$(f_{i},f_{j})=\frac{f_{i}}{f_{j}}$ (17)
A.3 Prompting Structure
At the beginning of the iteration, we provide the LLM with a prompt to guide it in generating features as expected. The prompt consists of three parts: (1) Feature Engineering, (2) Iterative Refinement and Evaluation, and (3) Markdown and Organization. Here are the detailed descriptions of these parts:
  1. 1.
    Feature Engineering. In the first round of iteration, $l$ agents independently employ the Tree of Thoughts method on the original feature set $\mathcal{F}$. They select, apply operations to, or delete features to generate new features and form new feature subsets $\mathcal{F}_{l}$.
  2. 2.
    Iterative Refinement and Evaluation. Then, we submit these $l$ feature subsets to downstream tasks to assess the performance of each $\mathcal{F}_{l}$, defined by the indicators $\mathbf{\theta_{1}}=(\theta^{1},\ldots,\theta^{l})$. We provide $\mathbf{\theta_{1}}$ and the performance of the original feature set $\mathbf{\theta_{0}}$ in the downstream tasks as feedback. Each expert shares their thought process in detail, including considering the ideas of others and acknowledging errors, and evaluates their performance based on the improvement in the downstream tasks.
  3. 3.
    Markdown and Organization. We organize the entire process in a Markdown table format, meticulously documenting every step of selection, operation, and deletion. As we enter the second iteration, each agent, based on their respective feature subset $\mathcal{F}_{l}$, generates new features guided by the evaluation and downstream task metrics.
In the prompt, we implement the Tree of Thoughts (ToT) methodology. Here, the process that each agent analyzes and selects features to operate on plays the role of Thought Decomposition. The key to this generation process is logical analysis and the prediction of potential impacts of features of the whole feature set. In this process, agents generate multiple potential solutions according to the current state of the feature set. At the end of each round, we evaluate the performance of new feature subsets in downstream tasks. This evaluation and the comparison of performance before and after play the role of the State Evaluator. This process provides a quantitative evaluation for the iterative optimization of the feature set. Through several iterations, the entire LLM continuously adjusts and optimizes, thus searching for the optimal feature set that can significantly enhance the performance of downstream tasks.

### v2

@@ 3.3 Feedback and Optimization
We now have a new feature set composed of all $l$ subsets, denoted as $\mathcal{F}_{l}^{{}^{\prime}}={\{\mathcal{F}_{l},g_{1},\ldots,g_{k}\}}$. After the agents generate all $l$ subsets, we move forward to evaluate the performance of these subsets in downstream tasks and the effectiveness of the generated nodes at this layer. We then gather performance metrics from downstream tasks as feedback $\mathbf{\theta}=\{\theta_{t}\}$, for each iteration $t$ corresponding to each generation layer. This feedback reflects effectiveness of the newly created feature subset and illustrates the agents’ contribution to the task. The agents evaluate their performance by comparing the metric $\theta_{t}$ with $\theta_{(t-1)}$. If $\theta_{t}$ significantly improves over $\theta_{(t-1)}$, the feature generation strategy is considered “effective”. The agents’ self-evaluation positively correlates with the increase in these metrics; a larger improvement signifies a more successful feature generation strategy.
Furthermore, each agent shares its thought process and inferences during generating features, including the logic behind selecting specific operations. This transparency enables agents to understand and adopt diverse strategies to achieve collaborative improvement. Through self-evaluation and peer interactions, agents identify which strategies successfully enhanced performance and which may require adjustments or replacements. In the subsequent rounds, agents generate increasingly effective feature subsets adopting $\mathcal{F}_{l}^{{}^{\prime}}$ (from the previous round) and $\mathcal{O}$ to form the next layer of generated nodes. The iterative process continues until the optimal feature set $\hat{\mathcal{F}}$ is found or the maximum number of iterations $T$ is reached:
@@
Monte Carlo Tree Search.
In order to further optimize the feature space and balance exploration with exploitation, we employ an improved Monte Carlo Tree Search (MCTS) to explore new or underutilized feature space structures after the initial $T$ iterations. We adopt this MCTS framework because of its random sampling-based search strategy, which enhances the probability of finding superior feature combinations. In our case, this flexible searching method constructs a tree from all agents’ generation nodes and adjusts the search direction based on the performance feedback. Here, we define the root node as the original feature set given to the LLM, with each generation node representing a tree node. The search covers all generation nodes in the last iteration layer $Layer_{T}$ as leaf nodes.
@@
We start with node evaluation by calculating the improvement in the downstream task’s performance compared to the parent node. The evaluation reflects the effectiveness of newly generated features. Then, the MCTS iteratively builds a search tree by cycling through four phases: selection, expansion, evaluation, and searching. In the selection phase, children nodes are recursively selected from the root utilizing the Upper Confidence Bound (UCB) Auer et al. (2002):
@@
where $s_{i}$ and $s^{{}^{\prime}}_{i}$ are the visit counts for node $i$ and its parent respectively, $C$ is a hyperparameter balancing exploration and exploitation. The value $w_{i}$ represents the average performance improvement of all descendant nodes of node $i$, computed based on changes in the downstream task performance metric $\mathbf{\theta}$. Specifically, for each node $i$, the value of $w_{i}$ is calculated as the average performance improvement observed from its child nodes in the downstream tasks. This is given by:
$w_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{j\in\mathcal{C}_{i}}(\theta_{t}^{(j)}-\theta_{(t-1)}^{(j)}).$ (6)
Here, $\mathcal{C}_{i}$ is the set of child nodes of node $i$, and $|\mathcal{C}_{i}|$ is its cardinality (i.e., the number of child nodes of $i$). $\Delta\theta^{(j)}=(\theta_{t}^{(j)}-\theta_{(t-1)}^{(j)})$ represents the performance improvement contributed by child node $j\in\mathcal{C}_{i}$, where $\theta_{t}^{(j)}$ and $\theta_{(t-1)}^{(j)}$ are the performance metrics associated with the state after exploring child $j$ and the state before, respectively.
In the tree, we select the most promising node $s_{l}$. We define the selection criterion for each node as:
@@
$\mathcal{F}_{l}^{\prime}=\mathcal{G}(\mathcal{F}_{l}^{\text{select}},\mathcal{O}^{\prime}),$ (8)
where $\mathcal{G}$ is the transformation function generating $\mathcal{F}_{l}^{{}^{\prime}}$ by applying operations from a selected subset $\mathcal{O}^{\prime}\subseteq\mathcal{O}$ (the predefined complete operation set) to $\mathcal{F}_{l}^{\text{select}}$. At the end of the MCTS process, the optimal feature subset is determined by choosing the leaf node with the highest $w_{i}$ value, representing the feature set that maximized performance improvement during simulations:
$\mathcal{F}^{*}=\mathcal{F}_{i^{*}},\quad\text{where}\quad i^{*}=\underset{i\in\text{LeafNodes}}{\arg\max}\;w_{i}.$ (9)
This $\mathcal{F}^{*}$ represents the optimal feature subset identified by the MCTS, which should then be validated further in real-world tasks to confirm its efficacy.
@@ 4 Experiments
In this section, we present three experiments to demonstrate the effectiveness of LFG. First, we compare LFG against several baseline methods on multiple downstream classification tasks. Second, we perform a robustness check on LFG’s performance improvement. Finally, we further study and analyze iterative performance improvements in experimental results and discuss their reasons.
@@ 4.1 Experimental Setup
Datasets.
We evaluate the LFG method on four real-world datasets from UCI, including Ionosphere (Ino Sigillito et al. (1989), Amazon Commerce Reviews (Ama) Liu (2011), and Abalone (Aba)  Nash et al. (1995), as well as Diabetes Health Indicators Dataset (Dia) Teboul (2022) from Kaggle. The detailed information is shown in Table 1. For each dataset, we randomly selected $55\%$ of data for training.
@@
Table 1: Datasets descriptions.
Metrics.
We evaluate the model performance by the following metrics: Overall Accuracy (Acc) measures the proportion of true results (both true positives and true negatives) in the total dataset. Precision (Prec) reflects the ratio of true positive predictions to all positive predictions for each class. Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class. F-Measure (F1) is the harmonic mean of precision and recall, providing a single score that balances both metrics.
Downstream Tasks.
We apply the LFG model across a range of classification models, including Random Forests (RF), Decision Tree (DT), K-Nearest Neighbor (KNN) and Multilayer Perceptrons (MLP). We compare the performance outcomes in these models both with and without our method.
Baseline Models.
We compare the LFG method with raw data (Raw), the Least Absolute Shrinkage and Selection Operator (Lasso), and Feature Engineering for Predictive Modeling using Reinforcement Learning Khurana et al. (2018) (RL). Here, we set the same operation set for both LFG and RL method to consist of square root, square, cosine, sine, tangent, exp, cube, log, reciprocal, sigmoid, plus, subtract, multiply, and divide. For our LLM, we perform all the experiments on OpenAI API, GPT-3.5 Turbo model OpenAI (2024).
@@ 4.2 Experimental Results
Overall Performance.
Table 2 shows the overall performance results for LFG and the baseline models.
(1) Comparison with baseline models. From the table we can see that the LFG method consistently surpasses baseline methods across a variety of metrics and datasets. Specifically, we show the results of LFG in 3 iterations of generation, denoted as LFG-3, and compare it with the full LFG in $T\leq 10$. On dataset Ion the LFG achieves the best accuracy of $95.6\%$, which is an $6.4\%$ increase over raw data on task RF, and $4.5\%$ over RL method. On dataset Ama, the LFG-3 improves the accuracy of raw data of $59.6\%$ to $61.5\%$ while LFG performs an increase of $4.3\%$ on RF. Beyond accuracy, LFG consistently demonstrates significant improvements over baseline methods across other metrics and datasets. For instance, on dataset Ion, the highest recall improvement of 9.8% on DT reflects LFG’s ability to effectively handle positive samples. Similarly, the F1 score improvements, such as increasing from 0.476 to 0.607 on dataset Ama on KNN and from 0.855 to 0.932 on dataset Ion on DT, showcase the model’s adaptability across diverse tasks and feature distributions.
(2) Performance across different models and metrics. In precision, the LFG also shows superior performance. For example, on dataset Ion the LFG reaches a precision of $90.6\%$, surpassing the highest precision of $86.7\%$ among all baselines on KNN. This result indicates that the LFG is effective in reducing misclassifications. Regarding the metric recall, the LFG increases the Ion’s recall by $9.8\%$ compared to vanilla data and by $4.3\%$ compared to best of baselines on DT task, which demonstrates that LFG brings higher improvement on identifying positive samples. For F1, LFG on Ama improves the data performance from $63.8\%$ to $67.6\%$ in three iterations, and the F1 score increases up to $68.0\%$. As F1 Score is the harmonic mean of precision and recall, this improvement shows the LFG’s capability to balance between reducing misclassifications and minimizing missed classifications.
@@
Robustness Check.
We validate the experiments through five-fold cross-validation. The LFG method demonstrates robust performance across various datasets and achieves notable metric improvements, including $2.72\%$ in accuracy and approximately $5\%$ in precision, recall and F1 scores with LFG-3, and further $4.81\%$ in accuracy and approximately $7\%$ in other metrics with LFG. The results illustrate the adaptability and effectiveness of the LFG. The extended iterations continue to refine and optimize the feature space, which is reflected in the continuous improvements in precision and recall. The consistent enhancement across all metrics demonstrates the LFG’s reliability and stability comprehensively.
Iterative Performance Improvements.
We then compare the performance improvements of LFG-3 and LFG. On one hand, the incremental improvement from LFG-3 to LFG illustrates the continuous enhancement in all metrics, indicating that during additional iterations, the model refines and adjusts its generation strategy to explore better solutions for the feature space. On the other hand, the results of LFG-3 with only three iterations already show notable enhancement in all datasets. This shows that the model effectively and rapidly finds the zones of optimal feature space within limited steps of searching. As shown in Figure 5, we can see that the average number of features in a feature subset increases continuously in the first $5$ rounds of generation. The increase in features implies that the feature generation algorithm is actively finding new, potential feature combinations in the feature space and can expand the feature space effectively.
@@ 5 Conclusion
In this paper, we present a novel feature generation method LFG utilizing LLMs to enhance automated feature engineering’s interpretability, adaptability, and strategic flexibility. Our target in designing the model is to address traditional feature generation challenges, including lack of explainability, limited applicability, and rigid strategy formulation. Our approach utilizes expert-level LLM agents to overcome these. These problems can significantly limit broader deployment of feature engineering on machine learning tasks in diverse scenarios. Thus, we present LFG with a dynamic, adaptive automated feature generation approach to enhance interpretability and extend utility across various data types and tasks. Moreover, we present extensive experimental results demonstrating our approach outperforms various existing methods. For future work, we plan to extend this automated feature engineering paradigm across various techniques and machine learning tasks, and investigate transformations our approach applies to the feature space to gain deeper insights into its underlying mechanisms and impacts.
@@ 6 Limitations and Ethics Statements
While our LFG shows significant advancements and wide adaptability, there are several limitations that require further exploration, including: (1) High computational demands and limited scalability with very large or complex datasets. With LLM agents being the generator, the computational demands are comparatively higher than traditional methods; (2) The effectiveness of the generated features heavily relies on the quality of input data, which can affect the model’s performance in scenarios with poorly curated or noisy datasets; (3) The current work focuses solely on the tabular feature generation process. More complex generation methods on different data are not considered. Extending the LFG to handle feature generation in non-tabular scenarios remains a challenge.
@@
The LFG framework can enhance the effectiveness and explainability of the feature generation process. However, as the framework uses pre-trained GPT-3.5 Turbo as the generation model, it may inherit the ethical concerns associated with GPT-3.5 Turbo, such as responding to harmful queries or exhibiting biased behaviors.
Acknowledgements
Pengfei Wang is supported by the National Natural Science Foundation of China (Grant Nos. 62406306 and 92470204).